If the tenth unit produced took 24 hours and the fifth unit took 32 hours, answer the following:

a) What is the learning rate?

b) How long did the first unit take to complete?

c) How long did the third unit to take to complete?
Answer #1

solution i riven if the tents out boduced took 24 (houts) and the fifth unit took 32 Chours). And the Given time taken for 207 (5) _7*(5). Dividing as T (20) and +(G) Here, 24/22 Giren by 20% cout=ayhr og and similarly 5th out = 32 hr 2/= 366) 1944)b) It takes 6240 hours to Produce first uut. c) Time (+3) = T (2 xb) = 6240* ( 20.0415) = 39.55 So, It takes as 39.55 hours t